OpenStudy (anonymous):

The base of a triangle is 8 cm greater than the height. The area is 42 cm^2. Find the height and length of the base.

OpenStudy (jlastino):

Let base = x+8 height = x Area of triangle (1/2)(base)(height) (1/2)(x+8)(x) (x^2+8x)/2= 42 x^2+8x=84 x^2+8x-84=0 (x-6)(x+14)=0 x=6 x=-14 You can't have negative measurements :| so x=6 x+8= 14 Base = 14 Height = 6